OpenStudy (anonymous):

Factor 25x2 + 10x + 1.

OpenStudy (anonymous):

\[25x^2 + 10x + 1 = 25x^2 + 5x +5x+1 = 5x(5x +1) + 1(5x+1)\] = \[(5x+1)(5x+1) = (5x+1)^2\]

OpenStudy (ganpat):

25x2 + 5x + 5x + 1 = 0 5x (5x + 1) +1 (5x + 1) = 0 (5x + 1) (5x + 1) = 0 (5x + 1)2 = 0 x= -(1/5)